# OneForecast: A Universal Framework for Global and Regional Weather Forecasting

### v2

@@ Appendix A Proofs of Theorems
Consider an improved multi-stream message passing mechanism. Let a graph signal $\mathbf{f}\in\mathbb{R}^{N}$ have a spectrum $\hat{\mathbf{f}}=\mathbf{U}^{\top}\mathbf{f}$ under the graph Fourier basis $\mathbf{U}=[\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{N}]$, where $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ is the normalized graph Laplacian and $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})$ with $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$ Define the frequency response function of the message passing operator by $\rho:\lambda\mapsto\mathbb{R}.$ If the dynamic gating weights satisfy
@@
Because $\lambda_{i}\in[0,2]$, when $\lambda_{i}$ is close to $2$, it represents a high-frequency component on the graph. In that regime, the gating weights become larger, and the message passing operator $\mathbf{G}$ amplifies those components. Similarly, when $\lambda_{i}$ is near $1$, the factor $|\lambda_{i}-1|$ can still be significant enough to enhance mid-to-high frequencies. Conversely, for $\lambda_{i}$ near 00 (low frequency), the gating is relatively small and thus tends to suppress those components.
@@ C.1 Dataset
Global Geopotential (Z) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Specific Humidity (Q) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Temperature (T) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
U Component of Wind (U) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
V Component of Wind (V) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
10 Metre U Wind Component (U10M) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
10 Metre V Wind Component (V10M) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
2 Metre Temperature (T2M) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Mean Sea Level Pressure (MSLP) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Regional Mean Sea Level Pressure (MSLP) 1 0.25° 6h 7.5°W114°Esimilar-$\sim$ 36°W172.5°E 1959$\sim$2020
10 Metre U Wind Component (U10M) 1 0.25° 6h 7.5°W114°Esimilar-$\sim$ 36°W172.5°E 1959$\sim$2020
@@ E.1 Earth-specific Region Refined Multi-scale Graph
Mesh Nodes. $\mathcal{V}$ is the ensemble of mesh nodes, which contains multi-scale mesh nodes of different fineness and region refined mesh nodes that cover the global area. The mesh nodes are distributed over a refined icosahedron that has undergone five levels of subdivision, and the coarsest icosahedron consists of 12 vertices and 20 triangular faces. By dividing each triangular face into four smaller triangles, an additional node is generated at the midpoint of each edge. The new nodes are then projected back onto the unit sphere, gradually refining the grid. To enhance the forecasting performance in key regions, we further refine specific areas of the finest mesh, achieving localized mesh densification. For the global forecast task, we refine the 2 areas: 0°N105°Esimilar-$\sim$30°N160°E and 10°N-95°Wsimilar-$\sim$30°N-35°W. The features of each node include the cosine value of the latitude, as well as the sine and cosine values of the longitude. We only keep the finest mesh nodes, since the nodes on the coarse mesh are its subset. In total, the graph structure of OneForecast comprises 12337 mesh nodes, each characterized by three features.
@@ E.2 Encoder
where, $\mathbf{W}_{e}$, $\mathbf{W}_{s}$, $\mathbf{W}_{d}$ are the linear transformation matrix of grid2mesh edge features, send node feature, and target node features. $\mathbf{W}_{\text{ESMLP }}$ is the linear transformation matrix of output layer. $b_{d}$ is the bias of mesh node during linear transformation, $b_{ESMLP}$ is the bias vector of ESMLP. In summary, the grid2mesh edge update process can be define as:
@@ F.3 Typhoon Tracking
This study focuses on two extreme cyclones: Tropical Storm Yagi and Severe Typhoon Molave. The Yagi formed near Ito To, Japan on August 6, 2018, and landed over Wenling, China on August 12. The Molave formed on October 11, 2020, and landed over the Philippines on October 25, 2020. The initial conditions for these two cyclones are set at 0:00 UTC, August 6, 2018 and 0:00 UTC, 11 October 2020, respectively. Since there is no Fuxi results in 2018 on WeatherBench2, we can not compare it with Yagi. The results of Best Track (Ying et al., 2014) (Lu et al., 2021) can be found in https://tcdata.typhoon.org.cn/en/.

### v3

@@ Appendix A Proofs of Theorems
Consider an improved multi-stream message passing mechanism. Let a graph signal $\mathbf{f}\in\mathbb{R}^{N}$ have a spectrum $\hat{\mathbf{f}}=\mathbf{U}^{\top}\mathbf{f}$ under the graph Fourier basis $\mathbf{U}=[\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{N}]$, where $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ is the normalized graph Laplacian and $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})$ with $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2.$ Define the frequency response function of the message passing operator by $\rho:\lambda\mapsto\mathbb{R}.$ If the dynamic gating weights satisfy
@@
Because $\lambda_{i}\in[0,2]$, when $\lambda_{i}$ is close to $2$, it represents a high-frequency component on the graph. In that regime, the gating weights become larger, and the message passing operator $\mathbf{G}$ amplifies those components. Similarly, when $\lambda_{i}$ is near $1$, the factor $|\lambda_{i}-1|$ can still be significant enough to enhance mid-to-high frequencies. Conversely, for $\lambda_{i}$ near 0 (low frequency), the gating is relatively small and thus tends to suppress those components.
@@ C.1 Dataset
[…]
@@ E.1 Earth-specific Region Refined Multi-scale Graph
[…]
@@ E.2 Encoder
[…]
@@ F.3 Typhoon Tracking
This study focuses on two extreme cyclones: Tropical Storm Yagi and Severe Typhoon Molave. The Yagi formed near Iwo Jima, Japan on August 6, 2018, and landed over Wenling, China on August 12. The Molave formed on October 11, 2020, and landed over the Philippines on October 25, 2020. The initial conditions for these two cyclones are set at 0:00 UTC, August 6, 2018 and 0:00 UTC, 11 October 2020, respectively. Since there is no Fuxi results in 2018 on WeatherBench2, we can not compare it with Yagi. The results of Best Track (Ying et al., 2014) (Lu et al., 2021) can be found in https://tcdata.typhoon.org.cn/en/.